What is the result of isolating x^2 in the equation? (x+1)^2+(y-8)^2=9
A.x^2 = -y^2 - 2x + 16y - 56
B.x^2=y^2+56
C.x^2=y^2+2x+16y+56
D.x^2=-y^2-56

Answer by Theo(13342)

start with (x+1)^2 + (y-8)^2 = 9

simplify to get x^2 + 2x + 1 + y^2 - 16y + 64 = 9

combine like terms to get x^2 + 2x + y^2 - 16y + 65 = 9

subtract 65 from both sides of the equation to get x^2 + 2x + y^2 - 16y = -56

subtract 2x from both sides of the equation and subtract y^2 from both sides of the equation and add 16y to both sides of the equation to get:

x^2 = -56 - 2x - y^2 + 16y

rearrange the terms in descending order of degree, with the x variable shown before the y variable when the degree is the same, to get:

x^2 = -y^2 - 2x + 16y - 56

that looks a lot like selection A, which is:

x^2 = -y^2 - 2x + 16y - 56